Suppose the central bank's monetary policy sets the interest rate according to the function:
i = 2 + 2(π - π*) with π* = 2,
and aggregate expenditure is:
AE = 140 + 0.2Y - 10i

a) What is the equation for the ADπ curve? Use 'PI' to represent π when entering your equation.

b) Plot the ADπ curve in the graph below.
1
Expert's answer
2020-11-19T08:27:14-0500

a)"AD=AE"

"i = 2 + 2(\u03c0 - \u03c0*)"

with π* = 2

"i = 2 + 2(\u03c0 -2)=2+2\u03c0-4=2\u03c0-2"

"AE = 140 + 0.2Y - 10i"

"i =2\u03c0-2"

"AE = 140 + 0.2Y - 10i=140+0.2Y-10(2\u03c0-2)=140+0.2Y-20\u03c0+20=160+0.2Y-20\u03c0"

"PI=\u03c0"

"AD=AE=Y"

"AD =160+0.2Y-20\u03c0"

"Y =160+0.2Y-20PI"

"Y=\frac{160-20PI}{0.8}"

"AD=\frac{160-20PI}{0.8}"

b)